### v1

# LLMs, You Can Evaluate It! Design of Multi-perspective Report Evaluation for Security Operation Centers

Abstract
Security operation centers (SOCs) often produce analysis reports on security incidents, and large language models (LLMs) will likely be used for this task in the near future. We postulate that a better understanding of how veteran analysts evaluate reports, including their feedback, can help produce analysis reports in SOCs. In this paper, we aim to leverage LLMs for analysis reports. To this end, we first construct a Analyst-wise checklist to reflect SOC practitioners’ opinions for analysis report evaluation through literature review and user study with SOC practitioners. Next, we design a novel LLM-based conceptual framework, named MESSALA, by further introducing two new techniques, granularization guideline and multi-perspective evaluation. MESSALA can maximize report evaluation and provide feedback on veteran SOC practitioners’ perceptions. When we conduct extensive experiments with MESSALA, the evaluation results by MESSALA are the closest to those of veteran SOC practitioners compared with the existing LLM-based methods. We then show two key insights. We also conduct qualitative analysis with MESSALA, and then identify that MESSALA can provide actionable items that are necessary for improving analysis reports.
1 Introduction
Cyberattacks have increased across many organizations, and establishing a security operation center (SOC) to analyze security alerts and provide responses has become urgent and crucial for each organization in recent years. An important mission for SOCs is, in addition to providing security alerts and their analyses, to write analysis reports with clear and actionable insights into their underlying cyberattacks [sharma2024decoding]. Interestingly, in proportion to the quality of analysis reports, various stakeholders, including not only analysts but also their managers, can understand the content of security alerts, and hence, the entire organization will be aware of the security [jawad2024m].
However, writing analysis reports for security alerts requires SOC practitioners to force a heavy workload, because these reports must contain actionable insights [kokulu2019MatchedandMismatchedSOCs]. To reduce this workload, methods to extract actionable insights for cybersecurity are in high demand [albanese2025towardsAI-driven]. Although there are several tools to generate analysis reports for security alerts [loumachi2025advancing, zhao2025information], automatically generated reports often contain inaccurate information, thereby affecting the SOC performance [kokulu2019MatchedandMismatchedSOCs]. Even when the state-of-the-art large language models (LLMs) are utilized, they cause hallucinations [CHEN2024SurveyoflargelanguagemodelsforCyberThreat, hasanov2024applicationoflargelanguagemodels]. Moreover, writing analysis reports to the exact requirements of report writing guidelines is stressful [nepal2024burnout]. The reason is that evaluation of analysis reports should identify any lack of descriptions as well as judgments from the perspectives of experts [RYAN2012Quantifying]: for example, “whether response is actionable from the content." The most common mitigation approach for the above problem is to constantly tune and adjust reports [kokulu2019MatchedandMismatchedSOCs], while the evaluation of analysis reports for SOCs needs expertise knowledge [kerstenfield2025tier1analyst].
Based on the above background, in this paper, we discuss the evaluation of analysis reports for security alerts in SOCs with LLMs in order to improve the quality of the reports. Specifically, we discuss the following research questions in this paper:
RQ1
What are the evaluation criteria that guarantee the quality of analysis reports from expertise knowledge?
RQ2
Can LLMs quantitatively evaluate the quality of analysis reports in accordance with judgment from the perspective of experts under the defined evaluation criteria?
RQ3
Can LLMs qualitatively evaluate the quality of analysis reports with feedback for judgment from the perspective of experts under the defined evaluation criteria?
We first propose Analyst-wise checklist as the evaluation criteria to reflect SOC practitioners’ knowledge for the analysis report as the answer to RQ1. To design this, we also conducted a literature review and a user study with semi-structured interviews. We collected 13 public guidelines and handbooks in IT and OT domains as thematic analysis in the literature review, and contacted 15 SOC practitioners with diverse backgrounds in these domains as template analysis [bushra202299False]. We also conduct a validation test of the checklist with multiple SOC practitioners. (See Section 3 for details.)
Next, under on the above checklist, we propose a novel framework, called Multi-perspective Evaluation System for Security Analysis using Llm Assistance (MESSALA), to maximize the evaluation of analysis reports in SOCs. In a nutshell, MESSALA guides LLMs in evaluating analysis reports by leveraging the Analyst-wise checklist from multiple perspectives. As described in Section 4, MESSALA can imitate human cognition to obtain expert knowledge by the checklist from the perspective of human experts, and therefore, can return appropriate evaluation results with feedback to the given analysis reports. When we conduct extensive experiments with MESSALA, we identify that MESSALA outperforms the existing methods [fu2023gptscore, liu2023g-eval] in quantitative evaluation. Thus, as the answer to RQ2, we confirm if MESSALA enables LLMs to quantitatively evaluate analysis reports from the perspective of human experts. (See Section 5 for details.)
Third, we examine whether the feedback returned from MESSALA aligns with judgments from the perspective of human experts through qualitative evaluation. According to the multi-metric rating, MESSALA obtains higher ratings in most metrics by virtue of feedback, which is more specific and easier to understand than common LLM baselines. It means that MESSALA also enables LLMs to qualitatively evaluate analysis reports through its feedback. (See Section 6 for details.)
To sum up, we make the following contributions in this paper:
  • We design an Analyst-wise checklist that contains items for analysis report evaluation in order to reflect SOC practitioners’ knowledge through literature review and semi-structured interviews as the answer to RQ1.
  • We propose MESSALA, a novel framework that maximizes report evaluation and provides feedback from the perspective of SOC practitioners, and confirm if MESSALA enables LLMs to quantitatively evaluate analysis reports through extensive experiments. This is identical to the answer to RQ2.
  • we examine the feedback returned from MESSALA, and then identify that MESSALA also enables LLMs to qualitatively evaluate analysis reports through its feedback, which is the answer to RQ3.
2 Related Work and Our Problem Setting
In this section, we describe related works in terms of LLMs and SOCs as background. We then describe evaluation of analysis reports in SOCs as the main problem setting.
2.0.1 Large-Language Models
LLMs are models that combine language processing with data generation. They learn observed tokens as sequences of characters and then find the most probable tokens in the output sequences [sandoval2023LostAtC]. Input to an LLM is called a prompt. LLMs can provide comparable performance to a veteran data analyst [cheng2023-IsGPTaGoodDataAnalyst], and are used to evaluate data as LLM-as-a-judge services [hu2024llm]. We focus on GPTScore [fu2023gptscore] and G-Eval [liu2023g-eval]. A common problem for LLMs is hallucinations with incorrect outputs [huang2025hallucination].
LLMs have been utilized in many cybersecurity research, such as data collection [BAYER2023multi-levelfine-tuning, papa2023AnAttackersDream, sharma2023HowWellDoesGPTPhishPeople], data pre-processing [liu2023NottheEndofStory, wadhwa2023RevisitingRelationExtraction], feature extraction [BOFFA2024Logprecis, jiang2024Xpert],attack detection [ali2023HuntGPT, Ghourabi2023EnhancingSpamMessage, koide2024ChatPhishDetector]and report generation [Gupta2023fromChatGPTtoThreatGPT, mitra2024LocalIntel, perrina2023agir, yamin2024applications]. However, LLMs require further fine-tuning for each specific task of cybersecurity [ferrag2023RecolutionizingCyberThreat]. We discuss a report evaluation method based on LLMs for cybersecurity.
2.0.2 Security Operation Centers
SOCs address various aspects of cyberattacks, including data collection, determining the presence of intrusions, and responding to threats [Tilbury2024HumansandAutomation]. The major role of SOCs is to detect cyberattacks [gupta2019AutomatedEventPrioritization, renners2019AdaptiveandIntelligible], investigate security alerts [yen2013Beehive, zhong2019LearningfromExpertsExperimence], and make judgments on these alerts [kokulu2019MatchedandMismatchedSOCs]. SOCs roughly consist of two processes, i.e., the evaluation process by analysts and the judgment process by managers. To support and connect both processes, creation of analysis reports is important [kokulu2019MatchedandMismatchedSOCs].
Although SOCs in recent years have introduced automation tools [AGYEPONG2023ASystematicMethod, shahjee2022IntegratedNetwork] to reduce workload [Tilbury2024HumansandAutomation], it is still heavy [nepal2024burnout, kerstenfield2025tier1analyst, kersten2024securityalert]. Analysts in SOCs also need to understand cyberattacks based on fragmented information [VANDERKLEIJ2022DevelopingDecisionSupport] and frequently require source information to determine appropriate responses [hassan2019NoDose, happa2021Assessing]. Although machine learning, including LLMs, has been developed for analysts in SOCs in recent years [ANDRADE2019CognitiveSecurity, ban2021CombatSecurityAlert, baroni2021Self-AwareEffective, gonzalez2021SIEM, HUSAK2022Crusoe, mitra2024LocalIntel, jiang2024Xpert, albanese2025towardsAI-driven, singh2025llms, cui2025trafficllm, li2024dollm, singh2025contextbuddy], there are limitations for providing analysis reports [michelet2024chatgpt] and security advice [chen2023can] from LLMs.
2.0.3 Evaluation of Analysis Reports in SOCs
We focus on evaluation of analysis reports in SOCs as the main problem setting. The current report generation tools have limitations since LLMs are often unable to generate analysis reports with high quality [michelet2024chatgpt]. Instead, we aim to explore whether LLMs can effectively evaluate these reports. We note that the evaluation of analysis report remains challenging: general LLMs may fail to align with human judgment because these reports often include domain-specific contexts [chiang2023closer].
Problem Setting
We aim to design a system that takes analysis reports as input and returns evaluation results as output. The output includes both quantitative scores and qualitative feedback from the perspectives of SOC analysts, such as actionable feedback. We consider that the readers of analysis reports are both analysts and managers, while the generation of the reports themselves are outside the scope of this paper.
3 Designing Analyst-wise Checklist for Security Report Evaluation
This section aims to clarify the criteria for evaluating the quality of security alert analysis reports and to construct the Analyst-wise checklist as a quality evaluation metric reflecting these criteria. The following section constitutes the Response to RQ1. This type of report is a highly specialized document that presupposes advanced security domain knowledge. It has been pointed out that general text evaluation metrics are insufficient for accurately assessing the quality of such domain-specific reports [stufflebeam2000guidelines]. The proposed Analyst-wise checklist integrates evaluation criteria that were iteratively validated through a literature review and interviews, and is designed to focus the assessment on key aspects that directly relate to practical decision-making.
3.1 Construction Process of the Analyst-wise checklist
The construction of the Analyst-wise checklist items was performed through the following phased and iterative process. This approach is a common practice for incorporating knowledge base utilization and user needs/field context into design [benton2020usability, kwong2025design, rose2022co]: 1) Collection of Candidate Items: Based on a review of security guidelines and existing related literature, we collected and formulated both the information that should be documented in analysis reports and the initial set of candidate evaluation items. 2) Interviews on Report Quality: Semi-structured interviews were conducted with SOC analysts to extract practical and critical evaluation perspectives on report quality. 3) Development of the Analyst-wise checklist Draft: We aligned the structural elements of analysis reports extracted from the literature review with the implicit evaluation criteria elicited from SOC analysts through interviews, and, based on these correspondences, newly designed concrete checklist items that are particularly important for assessing analysis report quality. 4) Refinement through Expert Review and Evaluation Tests: For the drafted checklist, we first revised the item structure and wording based on expert review. We then conducted trial evaluations in which multiple authors independently applied the checklist to sample reports, and, by reconciling major rating discrepancies and refining the phrasing of problematic items, confirmed an acceptable level of inter-rater agreement before finalizing the checklist. The details of this checklist are described below.
3.2 Collection of Candidate Items through Literature Review
The initial phase of developing the checklist items for analysis reports involved a systematic literature review. This review aimed to establish the design foundation of the checklist by ensuring the comprehensiveness and clarity of the information required for reporting alert analysis results. The process consisted of the following steps:
1) Literature Collection: To identify the knowledge required for the checklist items to possess pertinence, official guidelines concerning incident reporting and SOC operations, established by public organizations and expert communities, were set as the collection criteria. A total of 13 documents were targeted, including NIST SP 800-series incident handling guidelines and public information on incident reporting from domestic and international CSIRTs, regulatory bodies, and industry groups [johnson2003handbook, force2020security, johnson2016guide, scarfone2008sp, stouffer2023guide, brenner2007iso, cisa_incident_reporting_2024, enisa_eecc_incident_2021, uscert_incident_guidelines_2015, enisa_incident_management_2010, isogj_soc_textbook_2023, isogj_5w1h_2019, enisa_iot_sm_2019].
2) Extraction of Relevant Content: To ensure efficiency in the extraction process, we combined keyword searching, using terms such as "report," "should be described," "should be included," and "information sharing," with manual reading. This was done to extract sentences that enumerated the information required to be included in a report. For lengthy guidelines spanning dozens to hundreds of pages, we utilized an LLM. Specifically, we queried the LLM with "List the information that should be included in the incident report envisioned by this guideline, and specify where it is written." This provided a candidate list to efficiently narrow down the relevant sections.
3) Manual Selection of Statements: To ensure the clarity and concreteness required for checklist items, the statements obtained from keyword extraction and the LLM candidate lists were manually selected. By cross-referencing with the original text, only descriptions that could genuinely be considered "information items to be included in the report" were retained.
4) Organization of Information: The list of extracted passages was systematically coded by two of the authors with respect to the items that should be documented in an analysis report. Through this coding process, the extracted descriptions were conceptually grouped according to their functional role within the analysis report and organized into three overarching categories and eleven concrete documentation items, thereby establishing the foundation of the checklist. The draft set of three categories and eleven items was then reviewed by two SOC experts and subsequently revised. To ensure the comprehensiveness of the literature base, we extended the initial set of documents with several additional sources. As this extended review did not yield any new content that needed to be extracted, we regarded the data as saturated with respect to the information to be identified from the literature.
Below, we present the set of items that should be documented in an analysis report and the categories to which they belong.
Decision support and action planning: This category groups report elements that help frontline staff and managers grasp the current situation and determine appropriate next steps. The checklist items assess whether recommended actions and follow-up are concrete, executable, and well-justified, and whether they connect immediate response with longer-term improvement. [enisa_incident_management_2010, isogj_5w1h_2019, isogj_soc_textbook_2023, johnson2003handbook, force2020security, johnson2016guide, scarfone2008sp, uscert_incident_guidelines_2015]. Main report elements: Analysis Status; Impact Assessment; Confirmation Requests; Response Actions; Recurrence Prevention & Lessons.
Accountability and quality assurance of the investigation and analysis process: This category ensures that investigation activities and analytic judgments are documented in a transparent, traceable manner, supporting the reliability and auditability of SOC operations. The checklist items assess whether the process is described systematically and with sufficient detail to enable third-party review, compliance verification, and future incident handling. [stouffer2023guide, enisa_incident_management_2010, cisa_incident_reporting_2024, brenner2007iso, isogj_5w1h_2019, force2020security, johnson2016guide, scarfone2008sp, uscert_incident_guidelines_2015] Main report elements: Investigation & Analysis Methods; Evidence & Supporting Data; Related Policies, Guidelines & Standards.
Technical understanding and root cause clarification of the event: This category covers the technical description of what happened, where, to whom, and how. The checklist items assess whether the incident is described concretely enough to allow technical reconstruction and to justify the reported conclusions. [enisa_incident_management_2010, enisa_iot_sm_2019, enisa_eecc_incident_2021, johnson2003handbook, cisa_incident_reporting_2024, force2020security, scarfone2008sp, uscert_incident_guidelines_2015] Main report elements: Event Description & Interpretation; Root Cause; Incident Source & Impacted Systems; Communication Details; Vulnerability Information.
3.3 Interviews on Report Quality
Table 1: Participant information. For the row of Expertise, we categorize participants as “High", “Middle", and “Low" based on their self-reported expertise and years of experience.
User ID Expertise / Experience(years) Job Title Target Environment
1 High / 7 SOC Manager, Architect Factory, Building
2 High / 5 Engineer, SOC Analyst Building, Home Appliances
3 High / 9 SOC Manager, Architect IT, Factory
4 High / 6 SOC Analyst, Supervisor IT
5 Middle / 4 Engineer, SOC Analyst Factory, Home Appliances
6 High / 17 SOC Analyst Data Center
7 High / 23 SOC Analyst Data Center
8 Low / 3 SOC Analyst IT, Factory
9 High / 5 SOC Analyst, Supervisor IT, Factory
10 High / 12 SOC Analyst, Supervisor IT
11 High / 20 FSIRT/CSIRT Staff Factory
12 Middle / 2 PSIRT Staff Transportation, Cold Chain
13 High / 20 Control Vendor Security Staff Building
14 Low / 3 Security Manager (Manufacturing Site) IT, Factory
15 Low / 2 SOC Analyst, Network Engineer IT, Factory
We conducted semi-structured interviews with SOC personnel responsible for drafting and reviewing analysis reports in order to develop a checklist [rose2022co] that reflects SOC practitioners’ concerns about analysis reports and the operational context in which they are used. The purpose of these interviews was to elicit the implicit evaluation criteria that analysts rely on when judging the “quality” of a report and to make those criteria explicit as checklist items, focusing on the perspectives deemed particularly important. For the “report elements that should be documented” identified earlier, generic criteria and metrics for text quality can be applied to some extent. In this study, however, we deliberately focused on the perspectives that SOC practitioners themselves strongly recognize as problematic in their daily work, and prioritized clarifying which aspects should be checked most carefully. For example, while purely formal aspects such as spelling and typographical errors cannot be entirely ignored, it is more likely that, in analysis reports, practical “quality” is determined by content-related factors such as the technical accuracy of the detected event, the adequacy of the description of the operational context, and the concreteness and feasibility of the recommended response actions. We therefore considered it essential to capture these viewpoints appropriately and to ensure that they are explicitly and transparently represented in the checklist.
3.3.1 Study Design
We conducted interviews with 15 participants from 8 different SOC-related organizations. The participants included senior analysts responsible for alert analysis and customer-facing reports, team leaders, and managers, all of whom are routinely involved in at least one part of the report lifecycle—authoring, reviewing, or approving reports. Participants were recruited via a B2B interview platform, from within the authors’ own organization, and from partner companies. The participants included members of SOCs responsible not only for IT but also for OT/IoT environments, thereby ensuring diversity from multiple perspectives. Detailed demographic and organizational attributes of the participants are provided in the Table 1. All interviews were conducted online using conferencing tools such as Microsoft Teams. Wherever feasible, interviews were conducted by multiple authors to mitigate potential biases in questioning and interpretation. With participants’ consent, all interviews were audio-recorded and subsequently transcribed, yielding a total of 337 coded comments. Prior to analysis, the transcripts were anonymized and stored with appropriate security safeguards to ensure the ethical handling of the data. For participants from organizations external to the authors’ affiliation, we provided an honorarium set at a uniform amount across all participants. To allow the conversation to develop naturally while maintaining coverage of key topics, we adopted a semi-structured interview format, with each session lasting approximately 30 to 60 minutes.
As our analytic method, we employed template analysis (TA), which begins with a set of a priori themes of interest and iteratively refines and extends the coding template by incorporating newly emerging themes. TA is particularly useful in domains such as SOC research, where prior work is still limited and the researcher has only a partial understanding of the concepts that need to be identified in the data. [bushra202299False] First, we conducted an initial round of coding based on the descriptions obtained from the literature review and developed a preliminary set of thematic codes. We then analyzed the interview data, and whenever we encountered noteworthy segments that did not match any existing thematic code, we either introduced new codes or refined and extended the existing themes. Coding was performed iteratively three times, with intermediate code sets reviewed by SOC domain experts, and the final coding scheme was established based on the outcomes of these iterations. Moreover, as no new codes emerged after the 14th interview, we judged that thematic saturation had been reached and therefore did not conduct additional interviews beyond the 15 participants. In the following, we describe the evaluation criteria for assessing report quality that were derived from the interview results.
3.3.2 Implicit Evaluation Criteria for Security Report Quality
Table 2: Quality-oriented evaluation perspectives for analysis reports
Decision support and action planning
(1) Clarity and persuasiveness of the bottom-line conclusion for rapid situation understanding
(2) Quality of concreteness and prioritization in the proposed next actions
(3) Adequacy and clarity of the described customer or business impact
(4) Appropriateness of tailoring the content, tone, and level of detail to the recipient’s role and level of understanding
Technical understanding and root-cause clarification
(5) Accuracy and explanatory strength of factual descriptions, including explicit assumptions and constraints
(6) Appropriateness and depth of on-site contextual information usage, such as asset roles and operational background
(7) Adequacy of technical depth in describing the event beyond simple enumeration of logs
(8) Robustness of anomaly identification through the use of multiple analytical approaches, such as trends and comparisons
Accountability and quality assurance of the analysis process
(9) Effectiveness of structured presentation of key analytical elements in supporting accurate understanding of the analysis
(10) Clarity and verifiability of the analysis process through explicit and coherent links between evidence and conclusions
(11) Clarity of analysis scope and responsibility boundaries to facilitate seamless escalation and role demarcation
This section organizes the implicit criteria that SOC practitioners use when assessing the quality of security reports and refines them into explicit dimensions for evaluating report quality. Although most of the organizations we interviewed did not maintain formal documentation explicitly defining “report evaluation criteria,” we found that experienced Tier-2–level analysts, recipient-side managers, and SIRT members effectively applied similar criteria when reviewing reports. For an overview of interview findings other than the report evaluation criteria, refer to Table 2. Moreover, most comments pointed to shortcomings in the substantive content of the reports rather than in their formal qualities. In the following, we present the resulting evaluation criteria and related comments, organized according to the three categories of items that should be included in the report.
Evaluation perspectives for decision support and action planning
Evaluation criterion (1): Clarity and persuasiveness of the bottom-line conclusion for rapid situation understanding. Respondents emphasized that analysis reports should present a clear and persuasive bottom-line conclusion that enables recipients to rapidly grasp the situation and make informed decisions. In particular, reports are expected to explicitly state whether the alert is valid, its urgency, and the recommended stance toward the situation, together with the reasoning that supports these judgments. However, participants noted that many reports merely enumerate factual observations without articulating a decisive conclusion. “High-quality reports clearly state whether an alert should be treated as a true positive or not, how urgent it is, and why that judgment was made. Without that bottom line, it is hard to decide the next step quickly.” –- UserID 1. “For decision-making, the key is whether the report clearly explains what the problem is, why it matters, and what conclusion we should draw from the analysis.” –- UserID 3.
Evaluation criterion (2): Quality of concreteness and prioritization in the proposed next actions. Participants stressed that reports should not stop at analysis results but should concretely describe the next actions to be taken, along with their relative priority. Effective reports distinguish between actions that must be taken immediately and those that are optional or conditional, enabling recipients to allocate resources appropriately. In contrast, reports that simply list findings without actionable guidance were considered insufficient for operational use. “It is important to clearly state what should be done next and to indicate priorities, for example, whether an action is mandatory or should be handled if resources allow.” –- UserID 9. “Readers need to understand not only what issues exist, but also what actions they themselves are expected to take in response to those issues.” –- UserID 6.
Evaluation criterion (3): Adequacy and clarity of the described customer or business impact. Interviewees regarded it as essential that reports clearly describe how the alert affects the customer’s business, operations, or productivity, from the customer’s perspective. Reports that focus solely on technical details while leaving business impact ambiguous make it difficult for recipients to judge acceptability or risk. When possible, quantifying the impact was seen as particularly valuable for facilitating coordination and decision-making. “If the impact on business or operations is clearly described or even roughly quantified, it becomes much easier to explain the situation and move forward.” –- UserID 11. “Ultimately, customers want to know what caused the issue and whether the situation is truly acceptable. If that is unclear, the report leaves them uneasy.” –- UserID 14.
Evaluation criterion (4): Appropriateness of tailoring the content, tone, and level of detail to the recipient’s role and level of understanding. Respondents highlighted the importance of adjusting the content, tone, and level of technical detail according to the recipient’s role and expertise, in order to reduce cognitive burden and align perceptions of urgency. Participants pointed out that reports written in a uniform style often fail to bridge the gap between frontline analysts and decision-makers, resulting in misunderstandings or delayed responses. “Readability is critical. Each recipient has a different level of understanding, so the explanation and terminology need to be adjusted accordingly.” –- UserID 2. “What matters most is narrowing the gap in sense of urgency between the frontline and decision-makers. Reports should reflect the context and discussions that led up to the analysis.” –- UserID 10.
2) Evaluation perspectives for technical understanding and root cause clarification of the event
Evaluation criterion (5): Accuracy and explanatory strength of factual descriptions, including explicit assumptions and constraints Analysts emphasized that reports must accurately describe what occurred during the event and provide explanations that convincingly justify the conclusions drawn. Beyond factual correctness, respondents highlighted the importance of explicitly stating the assumptions, analytical grounds, and constraints underlying the analysis. Reports that omit these elements were considered difficult to trust, as readers cannot assess whether the conclusions are well-founded. “The most important thing is to accurately describe what actually happened and to explain it in a way that the reader can reasonably accept.” –- UserID 7. “If the assumptions, grounds, or constraints behind the analysis are unclear, it becomes hard to judge whether the situation has really been understood correctly.” –- UserID 4.
Evaluation criterion (6): Appropriateness and depth of on-site contextual information usage, such as asset roles and operational background. Participants reported that proper interpretation of alerts requires the use of on-site contextual information, including asset roles, system usage, and their position within operational or business processes. Without such context, analysts often cannot determine whether an alert represents a genuine issue or a false positive. Respondents noted that insufficient use of operational context frequently leads to overly conservative or ambiguous conclusions. “Operational context is indispensable. Depending on the role of the asset and how it is used, the priority and meaning of the alert can change completely.” –- UserID 13. “If information such as asset roles, whether the traffic is business-related, or results of on-site interviews is missing, we often cannot clearly determine whether an alert is a false positive.” –- UserID 5.
Evaluation criterion (7): Adequacy of technical depth in describing the event beyond simple enumeration of logs. Rather than merely listing communication records or log entries, reports are expected to provide technically meaningful explanations of the event, such as which systems communicated, for what purpose, and how that behavior should be interpreted. Several participants expressed frustration that the burden of interpreting raw technical data is often left entirely to the recipient. “SOC reports often just list IP addresses that communicated, and we have to re-investigate where those systems are and what they are used for.” –- UserID 6. “If the report also explained the meaning of the communication, it would significantly reduce the investigation workload on our side.” –- UserID 11.
Evaluation criterion (8): Robustness of anomaly identification through multiple analytical approaches, such as trends and comparisons. Finally, interviewees underscored the importance of identifying anomalies using multiple analytical perspectives, including temporal trends, comparisons with historical data, and correlations across different logs. Reports that rely solely on isolated log entries were considered insufficient for understanding whether observed behavior is truly abnormal. Participants noted that temporal and comparative analyses are still underutilized in many current reports. “What we really want to know is what has changed compared to the past, but that perspective is often missing in reports.” –- UserID 9. “Showing insights derived from differences in trends or combinations of logs would make anomalies much clearer.” –- UserID 3.
3) Evaluation perspectives for accountability and quality assurance of the investigation and analysis process
Evaluation criterion (9): Effectiveness of structured presentation of key analytical elements in supporting accurate understanding of the analysis. Participants emphasized that structuring and explicitly presenting key analytical elements—such as hypotheses, examined evidence, and interim judgments—is essential for ensuring that the reader can accurately understand the analysis. While standardized formats and procedures exist, respondents noted that differences in incidents and analyst experience often result in variation in how clearly these elements are articulated, leading to uneven comprehensibility across reports. “There is noticeable variation in report quality, and it often becomes dependent on the experience of individual analysts, especially in how clearly the analysis is structured.” –- UserID 4. “Even if formats are standardized, the way key analytical points are organized and explained still differs by analyst and by case.” –- UserID 1.
Evaluation criterion (10): Clarity and verifiability of the analysis process through explicit and coherent links between evidence and conclusions. It was considered critical that reports clearly document the reasoning process by explicitly linking evidence to conclusions, so that a third party could trace and verify how judgments were reached. Participants pointed out that reports often fail to sufficiently document how alternative explanations, such as false positives, were examined, which undermines confidence in the conclusions and the overall quality of the analysis. “It is important that there are no logical gaps or contradictions in how the analysis results lead to the conclusions, and that a third party could reproduce the same reasoning.” –- UserID 7. “Especially for determining whether an alert is a false positive, the evidence and reasoning must be clearly connected. If that part is weak, the quality of the entire report appears low.” –- UserID 11.
Evaluation criterion (11): Clarity of analysis scope and responsibility boundaries to facilitate seamless escalation and role demarcation. Participants emphasized that effective analysis reports clearly define the scope of investigation performed by the SOC and explicitly state responsibility boundaries for subsequent actions. High-quality reports do not attempt to exhaustively cover all aspects of an incident; instead, they clarify what has been analyzed within the SOC’s visibility, what conclusions can be reasonably drawn, and where responsibility is intentionally handed over to other stakeholders, such as CSIRT teams or customer-side organizations. This clarity facilitates seamless escalation, avoids redundant or speculative analysis beyond the analyst’s remit, and enables efficient collaboration across roles and organizational layers based on a shared understanding of ownership and accountability. “It is essential to clearly state what has been analyzed and what our assessment is, and then promptly escalate the case once it goes beyond our scope. Defining the boundary and handing it over early is crucial for enabling rapid and effective response.” –- UserID 12. “Because roles and scopes are clearly defined—both between SOC and CSIRT and across Tier 1, Tier 2, and senior analysts—we can escalate without hesitation and avoid over-analyzing matters that fall outside our responsibility.” –- UserID 11.
Table 3: Representative examples of categories, report elements, and checklist items with mapped evaluation criteria
Category / report element Example checklist item with mapped evaluation criteria
Decision support and action planning
Impact Assessment Are the on-site impacts, severity level, and risk evaluation described with justification? — (3)
Confirmation Requests Are specific items or questions to be confirmed by the recipient regarding this event described? — (2),(4)
Response Actions Are specific actions taken, their necessity, and the effects and evaluation methods described? — (1),(2)
Technical understanding and root cause clarification of the event
Event Description & Interpretation Does the report draw a concrete conclusion based on the analysis of the observed event? — (5),(7)
Root Cause If the incident is assumed to be an operational effect, is this explained? — (6),(7),(8)
Incident Source / Impacted Systems Is the origin device clearly identified along with its role, importance, and suspiciousness? — (6),(7)
Accountability and quality assurance of the investigation and analysis process
Investigation & Analysis Method Are the analysis methods, decision criteria, and viewpoints explained in a multi-faceted manner? — (9)
Evidence & Supporting Data Is evidence provided to support each observation or response, with source and acquisition method? — (10)
Analysis Status Is the current analysis progress and the next steps clearly stated? —(11)
3.4 Development of the Analyst-wise Checklist Draft
This checklist was developed by systematically integrating 1) the structural elements that should be included in analysis reports, extracted through a literature review, with 2) the implicit evaluation criteria used in practice by SOC analysts, elicited through interviews. The integrated item set was then finalized through expert review and application tests. This approach is also widely known as a general method for constructing evaluation checklists.[hales2008development]
Integration Method
To construct the checklist items, we first mapped, for each structural element of the analysis report, the evaluation perspectives derived from practitioners’ perceived challenges summarized in Section 3.3.2. We then defined concrete evaluation items by drawing on interview comments and descriptions of best practices extracted from the literature review.
When designing checklist items, our primary requirement is that they remain actionable in day-to-day SOC operations. Simply turning interviewees’ concerns into items risks creating demands that exceed the authority or resources of a typical SOC organization[rose2022co]. We therefore introduced a step that distills practically feasible elements from the interview findings and refines them into checklist items. For example, for evaluation criterion (6): use of on-site contextual information, interviewees repeatedly reported that the lack of local context often prevents them from deciding whether an alert is a false positive. However, it is unrealistic for a standard SOC to maintain detailed, up-to-date business context for every customer environment. Accordingly, our checklist does not merely ask whether incidents are analyzed with customer context in mind, but also whether analysts explicitly state which contextual information was available and organize missing yet decision-critical information as concrete confirmation requests. This yields a realistic checklist that assumes a division of roles with on-site staff while still reducing ambiguity caused by insufficient context.
The Analyst-wise checklist is also applicable to general tasks by human in SOCs because of the following concept: 1) the granularity of checklist items are generalized for the use in multiple domains, 2) for operational difficulty and flexibility in real-world systems, specific evaluation criteria are removed, 3) instead of applying the entire checklist, only relevant items for each analysis report are chosen.
These integration steps were primarily conducted by two researchers, including the first author. In addition, we used a general-purpose LLM (gpt-4.1) as a brainstorming aid: given the organized evaluation perspectives and structural elements, we prompted the model to “enumerate candidate checklist items for each perspective,” and used the generated suggestions to expand and refine candidate items and their wording. Before the expert review, the draft items were discussed among the authors and refined to retain only promising candidates.
3.5 Refinement through Expert Review and Evaluation Tests
The effectiveness of the constructed checklist was examined through expert review and a small-scale application test. First, three SOC domain experts reviewed the checklist and provided feedback on whether they would be willing to use it in practice, whether it helps prevent omissions during review, and whether its contents are consistent with the activities that should be performed in a SOC. Specifically, we received comments grounded in actual SOC practice, such as: “It is not the SOC’s role to prescribe concrete response actions, but they are expected to provide their assessment; therefore, such strictness should not be directly reflected in the checklist items.” Based on these results, we added, merged, and refined the wording of several items, conducted a second round of review, and, with reference to the Delphi method [rose2022co], selected as consensus items those that at least two of the three experts evaluated as useful.
As a final validation step, one of the three experts and one of the authors with more than seven years of SOC-related experience independently applied the checklist to ten real-world security alert analysis reports. For each checklist item, they rated the extent to which it was satisfied on a five-point scale and recorded the results. For items where their ratings diverged substantially in direction, e.g., one judged the item to be largely satisfied while the other judged it to be hardly satisfied, they reconciled their interpretations and refined the item descriptions and wording accordingly.
Ultimately, we confirmed that for more than 60% of the checklist items, the two raters agreed on the direction of judgment, i.e., whether the item was satisfied or not, and we then finalized the checklist. Table 3 shows a subset of the checklist; the complete checklist is provided in the appendix.
Ethical considerations regarding the use of interview results are described in Section 7.0.2.
3.6 Answer to RQ1
Based on the findings from the literature review and the semi-structured interviews, we define the quality of an analysis report as the following three standpoints: 1) Decision support and action planning for stakeholders, determining what actions are necessary and why; 2) Technical understanding and cause clarification by analysts, providing accurate descriptions of system behavior and its operations; 3) Accountability and reproducibility of the analysis process for each organization, ensuring that conclusions of the report are transparently linked to its evidence and reasoning. The quality of analysis reports for SOCs is judged by evaluation criteria from multiple perspectives of experts. Our Analyst-wise checklist contains these evaluation criteria.
4 Proposed Method: MESSALA
In this section, we propose MESSALA to overcome the challenges in providing feedback to analysis report creators using the Analyst-wise checklist constructed in the previous section. We first explain the main ideas behind MESSALA and then introduce the details of MESSALA, including its overall workflow.
4.1 Motivation and Design Requirements
To address the challenges of providing appropriate feedback to report authors, we introduce the following key idea. First, as discussed in in section 1 and  2, to overcome the challenges of dependence on veteran analysts, limited reviewer time and headcount, and the need for rapid feedback, we introduce an LLM-based automatic evaluation mechanism. The Analyst-wise checklist developed in the previous section is used, together with the report, as part of the LLM prompt so that the model can perform an evaluation from an expert-oriented perspective.
However, prior work has pointed out that LLM-based judgments of whether a text is good or bad tend to diverge from expert evaluations.[murugadoss2025evaluating] We also empirically examined this issue: we first had an LLM naively evaluate multiple sample reports using our Analyst-wise checklist, and then, on a later occasion, asked 5 of the 15 veteran analysts who participated in the interviews, those who provided consent, to review these LLM-based evaluations and provide comments. The results were consistent with the findings of prior work [murugadoss2025evaluating]. See the appendixXXX for the comments. 1) The tendency of the LLM’s quality judgments, good vs. bad, differed from that of the human analysts. 2) The feedback comments generated by the LLM were not effective. To provide appropriate feedback, an LLM must judge whether a report is good or bad in a way that is consistent with expert assessments and accurately synthesize these judgments into explicit feedback comments [zhou2024llm]. In this paper, within the context of security analysis reports, we formulate these two requirements as RQ2 and RQ3. As our response to RQ2 and RQ3, we propose MESSALA, a method that transforms the LLM-based evaluation process for analysis reports. MESSALA incorporates the following key ideas.
4.2 Key Idea
MESSALA rests on two key ideas. First, the Granularization Guideline steers the LLM’s reasoning with analyst expertise and practical feedback by translating the Analyst-wise checklist into fine-grained, actionable evaluation cues. This concentrates the model’s attention on expert-relevant aspects of a report.
Furthermore, to better approximate the quality of senior analysts’ judgments of report quality and their effective feedback, we designed a new architecture, the Multi-perspective Evaluation LLM. This architecture integrates a two-stage evaluation process. 1) High-level Evaluation: An initial assessment conducted based on simple prompt configuration and the surface information of the report. 2) In-depth Evaluation: A detailed assessment performed by deeply considering the report’s context through the application of the Granularization Guideline. The integration of results from both evaluations enables a multi-perspective evaluation. This design mimics the cognitive model of human text comprehension [wharton1991overview]. Humans first activate knowledge fragments through surface features during a construction phase, and then integrate their context during an integration phase to understand the overall meaning of the text.
By reproducing this human cognitive process, the Multi-perspective Evaluation LLM allows the LLM to more deeply understand the content of the Analyst-wise checklist, which was constructed from an expert perspective. Consequently, MESSALA is expected to achieve the evaluation and feedback provision of analysis reports at a level comparable to that of veteran analysts.
4.3 Detail of the Method
Figure 1: Overview of MESSALA. The top module runs two parallel evaluators: a High-level LLM and an In-depth LLM guided by the Granularization Guideline; their outputs are fused by the Multi-perspective Evaluation LLM.
The overall framework of MESSALA is illustrated in Fig. 1. MESSALA consists of three components: Analyst-wise checklist, granularization guideline, and multi-perspective evaluation. Since the Analyst-wise checklist was explained in the previous section, the following will describe the remaining two constituent components.
4.3.1 Granularization Guideline
The granularization guideline is defined with each category of the Analyst-wise checklist. It specifies, for instance, what types of contexts should be provided to an LLM, and what kind of descriptions should be emphasized during evaluation. This guideline functions as an abstract how-to-check rule that guides the LLM’s evaluation judgments, whereas the Analyst-wise checklist specifies what-to-check. The granularization guideline for each category is shown in Table 4. For example, in Category 1: “Hypothesis Validation," it focuses on guiding the LLM to find evidence for the assumption within the report.
The above categories of the granularization guideline are designed to evaluate whether actual analysis methods utilized in SOC are appropriately reflected to analysis reports, and whether sufficient contexts for the analysis is presented. They also evaluate whether hypotheses are explained with reasonable justification as shown in Table 4. The categories were iteratively refined through evaluation of analysis reports for SOCs in the real worlds. Furthermore, revisions were made after checking the categories with two SOC-related experts affiliated with the authors’ organization.
The granularization guideline prompts mainly address the following points: “What is the objective of this checklist item?"; “Which parts or types of information in the report should be examined?"; and “How should the identified information be evaluated?". Examples of the prompts used for each guideline category are shown in Fig. 6 in Appendix 0.C. The design of this guideline follows existing field studies for supporting analysis in SOCs [kersten2023give, kerstenfield2025tier1analyst].
By leveraging the granularization guideline, the in-depth evaluation LLM can be guided for detailed evaluation. It enables LLMs to evaluate not only the presence of content, but also consistency for contexts of analysis reports. The granularization guideline plays a crucial role in the in-depth evaluation LLM, as illustrated in Fig. 2. The granularization utilizes it to gain a deeper understanding for each item in the Analyst-wise checklist. It then determine which parts of the given analysis report is important. Finally, the in-depth evaluation LLM returns a part of input for the multi-perspective evaluation LLM described later.
Figure 2: Granularization guideline. The granularization guideline can break down items in the checklist into specific evaluation steps for each category.
Table 4: Overview of evaluation categories in the granularization guideline.
Cat. Name Detail
1 Hypothesis Validation Checks if the report provides sufficient reasoning and evidence for key conclusions.
2 Basic Alert Information Checks if key alert details (time, location, affected devices) are included.
3 Pattern and Comparison Analysis Checks if the report compares the incident with past cases or known patterns.
4 Detailed Communication Analysis Checks for communication analysis with technical context.
5 Surrounding Context and Timing Evaluates analysis of surrounding and time-related communications involving the alert device.
4.3.2 Multi-perspective Evaluation LLM with High-level and In-depth Evaluations
Our multi-perspective evaluation LLM unifies two evaluations, i.e., high-level evaluation and in-depth evaluation by LLMs. We first describe each evaluation below, and then describe the multi-perspective evaluation.
High-level Evaluation LLM
The high-level evaluation allows an LLM to rely on simple evaluation items with superficial information, such as grammatical correctness. For such an evaluation, the correlation between LLM outputs and human ratings is often high [murugadoss2025evaluating], and the use of expert knowledge may negatively affect with quality of evaluation. Meanwhile, when LLMs perform document evaluation without detailed guideline, they often produce holistic evaluations [zhou2024llm]. This may lead to insufficient attention, resulting in different analysis from veteran analysts. The prompt used for high-level evaluation is shown in Fig. 7 in Appendix. It consists of three components: the target security report, specific items in the Analyst-wise checklist, and the evaluation setup, including the evaluation task and its method. The evaluation is performed using a five-point Likert scale, which can provide human-like analysis [parker2024large, murugadoss2025evaluating] as well as the reliability of a rating scheme [kusmaryono2022number]. We note that, within the prompt, only a minimal evaluation method is provided by following the prompt structures in prior study [fu2023gptscore].
In-depth Evaluation LLM
The in-depth evaluation guides the LLM in understanding which descriptions are important for each item. This enables the LLM to perform evaluations based on contexts of analysis reports and approximates judgement from the standpoints of veteran analysts. The in-depth evaluation offers several advantages. First, it enables the prompt to provide the LLM with richer contexts, thereby enabling for rigorous evaluations. Second, it improves the interpretation of the evaluation process, and makes the LLM easier to justify the output. It also evaluate the granularization for each item automatically by LLMs [liu2023g-eval, lee2024checkeval]. We also note that evaluation based on general LLMs may fail to align with human judgment due to analysis reports, which include domain-specific contexts and differ from general documents [chiang2023closer], and prompts containing detailed instructions are sometimes ineffective [murugadoss2025evaluating] as described in the previous subsection. To overcome the above problem, The prompt used for the in-depth evaluation consists of three key components: an analysis report to be evaluated, the Analyst-wise checklist, and the granularization guideline as shown in Fig. 7 in Appendix. The task of the LLM is to generate a granularized version of evaluation based on content of analysis reports.
Multi-perspective Evaluation LLM
This step integrates high-level and in-depth evaluations in order to evaluate analysis reports from a perspective closer to veteran analysts. Since each evaluation method has distinct characteristics, leveraging their outputs is expected to yield more reliable and interpretable evaluation results. Moreover, by cross-referencing the outputs obtained from different levels of evaluations, it helps to mitigate both potential blind spots and biased judgments that arise from relying on a single viewpoint.
The prompt is structured as follows. First, based on the in-depth evaluation, the LLM is instructed to generate an initial score for the items in the multi-perspective evaluation checklist, which consists of the outputs of both evaluations. After reconsidering the initial scores, the LLM outputs a final score with a five-point Lkert scale. The prompt example is shown in Fig. 7 in Appendix.
5 Quantitative Evaluation of LLM Scores
In this section, we conduct extensive experiments for quantitative evaluations to determine whether MESSALA can evaluate analysis reports consistently with human experts, thereby addressing RQ2. Specifically, we compare the 5-point Likert scales evaluated by the LLM with those by human experts for analysis reports collected from real-world SOCs. We first outline the experimental setup, and then present the results.
5.1 Experimental Setting
We describe datasets, evaluation metrics, and baselines, including their implementations.
5.1.1 Datasets
We used the following two datasets for this experiments.
Real-world Analysis Reports
The first dataset is a private dataset as a collection of analysis reports produced by real-world SOCs. These reports were gathered from multiple SOCs monitoring three distinct environments, including factories, buildings, and IT infrastructure. Each SOC has independent operation and environment and differs substantially in its monitoring scope, IDS configuration, and alert types. All the analysis reports contain alerts triggered by network monitoring, although we omit their details due to ethical reasons as described in Section 7.0.2.
The dataset consists of 40 reports collected between April 2022 and April 2024, where 12 are for factories, 18 are for buildings, and 10 are for IT infrastructure. Report formats are independently standardized within each environment, and all the reports conclude that the alert represents a benign case; however, each report describes either events with non-negligible risk or ambiguous behavior, which require confirmation of the respective clients. Their contents include basic metadata (e.g., date and alert information), an event summary, impact assessment, detailed analysis, items requiring confirmation and action, and recurrence prevention. The average length of the reports is about 2,200 characters, and most reports are in PDF format. Sensitive information, such as IP addresses and proper names, was anonymized for the use of LLMs in a local environment. We do not plan to release the reports themselves due to ethical reason.
pseudo analysis reports
The second dataset is a pseudo analysis report dataset generated by an LLM. We use this dataset because, at writing this paper, there is no public dataset of analysis reports. To guarantee the reproducibility of this evaluation, we adopt an LLM-based pseudo–data generation approach [pmlr-v252-hao24a], which is commonly used in domains where real-world data are difficult to release. In particular, we generated pseudo analysis reports using a RAG-augmented CoT pipeline that incorporated MITRE ATT&CK  [^1] https://attack.mitre.org/ patterns.Through validating these pseudo analysis reports, 10 reports were selected for the final dataset.
For pseudo analysis report generation, we first prepared a report skeleton and CoT style prompt. The skeleton defined a basic structure in which environmental context and alert overview, analysis status, analytical conclusion, recommended actions, and supporting evidence are described in a consistent order. The CoT prompt specified a high-level generation procedure to ensure that these elements are described coherently. Because the skeleton and procedure alone would result in overly abstract content, we injected external knowledge and experiential information at the early stage of generation to increase concreteness. Concretely, we provided the LLM with (i) attack patterns from MITRE ATT&CK, (ii) typical patterns of detection, conclusion, and response extracted from past internal reports, and (iii) knowledge about hypothetical critical assets. In addition, for each target attack technique, we had the LLM refer to external technical reports so that the generated analyses and countermeasures would better reflect environment- and situation-specific interpretations. These steps were implemented by combining prompt chaining with a RAG-style retrieval process. After generating a large number of pseudo analysis reports through this pipeline, we performed automatic quality scoring using an LLM and iteratively discarded low-quality samples. Finally, we conducted a manual quality review and selected 10 pseudo analysis reports, which we used as the second dataset in our evaluation experiments.
Example prompts and representative generated pseudo analysis reports are provided in Appendix 0.F.
Human Gold Evaluation of Analyst-wise checklist
Using the Analyst-wise checklist in Section 3, we were also able to obtain the evaluation results from the five participant in the user study in Section 3. Each item in the Analyst-wise checklist was rated on a five-point Likert scale, and the final score was computed as the average of their individual scores. To align our understanding of the evaluation criteria, the first 20 reports were jointly reviewed. The remaining 30 reports were then processed by the two authors. For this human gold evaluation, 10 to 15 checklist items were selected, focusing on important contexts of the analysis reports. This limitation in the number of the items is based on two reasons: 1) Evaluating too many items would increase the workload on the participants and may cause low accuracy; 2) In realistic operational settings, it is more practical to concentrate on key items aligned with the content of the analysis reports in order to provide meaningful feedback.
In constructing the expert evaluation dataset, we followed three important precautions: 1) To minimize bias, evaluators were assigned only to reports from SOC environments they were not directly responsible for. As a result, each report was typically scored by approximately three evaluators, and their average was used as the final expert rating. 2) To ensure consistency among evaluators, we conducted preliminary briefings on the checklist items and held a Q&A session using Microsoft Teams to align understanding. 3) To preserve independence during the evaluation process, evaluators were not allowed to view each other scores until all evaluations were completed.
5.1.2 Evaluation Metrics
Following the discussion in Section 4.3.2 and a typical setting in LLM-as-a-judge [lee2024checkeval, liu2023g-eval, fu2023gptscore], we measure a five-point Likert scale and compute its correlation coefficients between an LLM and human as the primary evaluation metric. We compute the Spearman’s rank correlation $\rho$, the Kendall’s tau $\tau$, and the Pearson’s correlation $r$ as correlation coefficients, as well as the root mean square error (RMSE) to measure deviation in scores. While prior work [fu2023gptscore] computes and averages correlations for each individual document, we aggregate all analysis reports and then compute the correlations because the number of checklist items is limited and distinct for each report. We use a total of 755 items derived from 50 analysis reports on $n$ times execution.
5.1.3 Baseline
We implemented MESSALA and the following four methods as baselines including also ablation study, i.e., Method 1, Method 2, Method 3, and Method 4. Each method is evaluated using the same set of multiple models. Hyperparameters are temperature =0=0, top-$p=0$, and $n=5$. All methods use the Analyst-wise checklist.
Method 1: Only High-level Evaluation. This method performs a high-level evaluation described in Section 4.3.2. It follows GPTScore [fu2023gptscore] as a typical prompt that simply evaluates the given texts, leveraging the analyst-wise checklist while not applying the Granularization Guideline, and relying on a high-level evaluation prompt.
Method 2: Only In-depth Evaluation without Granularization Guideline. This method follows the G-Eval framework [liu2023g-eval] by first granularizing each item of the Analyst-wise checklist into a set of more in-depth evaluation criteria and then conducting the evaluation based on these detailed criteria. Notably, while this process introduces a granularization step, it does not employ our Granularization Guideline.
Method 3: Only In-depth Evaluation with Granularization Guideline This method performs the in-depth evaluation using both the Analyst-wise checklist and the Granularization Guideline compared with the previous methods. Note that it incorporates neither the high-level evaluation nor the multi-perspective evaluation.
Method 4: Multi-perspective Evaluation without Granularization Guideline In this method, the high-level evaluation results produced by Method 1 and the results of granularized items obtained in a manner similar to Method 2, where the Granularization Guideline is not applied, are provided as inputs to an LLM, which then outputs an integrated evaluation result.
5.2 Results
We present the results below, summarizing the numerical scores of each method and comparing MESSALA with the baseline approaches to answer RQ2. An overview of the results is shown in Table 5 and Fig. 3. As shown in the table, MESSALA outperforms all other methods across all metrics in most cases. These results highlight the advantages obtained from our two key modules, i.e., 1) the Analyst-wise checklist and the Granularization Guidelines, and 2) the multi-perspective evaluation with high-level and in-depth evaluations. Likewise, Fig. 3 shows that MESSALA exhibits the distribution closest to the human gold evaluation. These results demonstrate that MESSALA is significantly effective for evaluations of analysis reports because it is more aligned with human judgment than the other methods.
We also found several insights in an ablation study. When combining high-level and detailed evaluations, MESSALA consistently shows strong alignment with the human gold evaluation by grounding its judgments based on concrete report content and the Analyst-wise checklist instead of relying on superficial justifications. In contrast, using Method 1 alone tends to yield higher evaluation scores, as it relies on coarse, high-level assessments and lacks the ability to critically evaluate the detailed content of reports.
When comparing Method 2 and Method 3, which rely solely on in-depth evaluations, we observe distinct failure patterns, with Method 2 exhibiting inconsistent evaluation behavior due to the absence of Granularization Guidelines. As a result, Method 2 sometimes fails to specify evaluation points in sufficient detail, leading to overly generic assessments. Moreover, we observed multiple cases where Method 2 conducted evaluations that were weakly related or unrelated to the report content, resulting in low scores. In contrast, Method 3 consistently assigns excessively low scores even when only minor issues are identified. This tendency stems from its fine-grained and localized inspection process, which encourages stricter judgments, as also reported in prior work [kim2023prometheus].
Finally, a comparison between Method 4 and MESSALA shows that both approaches consistently produce results close to human evaluations across different models, demonstrating the benefit of multi-perspective evaluation. Furthermore, MESSALA outperforms Method 4, indicating that simply aggregating multiple evaluations is insufficient. These results highlight that stable and practical evaluation of analysis reports requires the integration of high-level and in-depth assessments guided by Granularization Guideline.
Table 5: Comparative analysis of closed models (left) and open models (right).
(a) closed models
Model $\rho$ $\tau$ $r$ RMSE
gpt-4o Method 1 0.54 0.46 0.54 1.16
Method 2 0.51 0.42 0.51 1.05
Method 3 0.49 0.40 0.49 1.27
Method 4 0.55 0.45 0.53 1.01
MESSALA 0.58 0.49 0.58 1.04
gpt-4.1 Method 1 0.63 0.53 0.61 1.27
Method 2 0.60 0.50 0.57 1.07
Method 3 0.49 0.40 0.49 1.30
Method 4 0.60 0.50 0.60 0.97
MESSALA 0.64 0.53 0.64 0.88
(b) open models
Model $\rho$ $\tau$ $r$ RMSE
gpt-oss (20B) 0.56 0.44 0.55 1.27
0.41 0.33 0.39 1.56
0.56 0.43 0.54 1.16
0.46 0.36 0.44 1.34
0.59 0.46 0.59 1.13
qwen3 (14B) 0.56 0.44 0.55 1.01
0.52 0.40 0.51 0.97
0.50 0.39 0.49 1.00
0.57 0.44 0.56 0.93
0.58 0.46 0.56 0.97
(a) Violin plot of method-wise scores from GPT-4.1.
(b) Violin plot of method-wise scores from GPT-4.0.
Figure 3: Violin plots illustrating the distribution of method-wise evaluation scores produced by each model, compared with human gold standard evaluations.
5.3 Answers to RQ2
For all the metrics, MESSALA consistently outperforms all the baseline methods: in particular, it achieves the highest correlation with the human gold evaluations while exhibiting the lowest score deviation and distributional divergence. These results indicate that an LLM can quantitatively evaluate analysis reports aligned with expert scores by virtue of MESSALA as the answer to RQ2.
6 Qualitative Evaluation of LLM Feedback Comments
In this section, we conduct experiments for qualitative evaluation in terms of the feedback comments generated by MESSALA. The goal of this evaluation is to identify whether the LLM can produce actionable and meaningful feedback from perspective of SOC practitioners. This experiment addresses RQ3. To this end, we design two complementary evaluations. The first evaluation examines how the LLM-generated feedback comments to analysis reports are useful for human experts and an LLM from qualitative standpoints, following prior work [zubaer2025gpt]. The second evaluation examines whether an LLM can correctly point out defects within analysis reports as feedback using a dataset consisting of defect-injected analysis reports.
6.1 Multi-metric Rating by Human Experts and an LLM Judge
We describe Multi-metric Rating evaluation assessed by both human analysts and an LLM as a qualitative evaluation of the feedback comments generated by the LLM.
6.1.1 Experimental Setting
3 participants (UserID 1, 2, and 5), who also took part in the interviews described in Section 3, were involved in this evaluation. They are veteran analysts and managers in SOC of the authors’ affiliation and belong to a different SOC domain, such as factories, buildings, and home appliances. Each participant was assigned 6 distinct analysis reports, consisting of two reports from each of the three domains, thereby evaluating feedback comments on 18 analysis reports. The task for each report was limited within approximately 10–15 minutes, and the total task per participant was completed around one hour to reduce their workload. We also applied the same evaluation to an LLM for the entire dataset, which is a common approach for assisting human evaluation [zubaer2025gpt]. This additional evaluation enables us to obtain more stable results with respect to the results in human analysts.
Considering the interviews in Section 3, we adopt the following metrics: Overall Usefulness to measure how feedback is practical; Accuracy and Validity to measure how feedback is precise; Specificity and Concreteness to measure how feedback is actionable; and Support for Novices/Veterans to measure how feedback supports analysts with different experience levels from novices to veterans. We conducted a user questionnaire on the above metrics using a 5-point Likert scale, where “1" indicates “very low" and “5" indicates “very high". Participants can also provide free-text comments to explain their ratings. We also utilize Method 1 described in the previous section, where LLMs are GPT-4o and GPT-4.1, as a baseline, and compute the average of their results.
Feedback comments by LLM
We limit the length of feedback comments to 500 tokens by summarizing the results derived from each method. While MESSALA sometimes generates feedback comments with excessive information as noted earlier, a suitable amount of feedback for users is around 500-700 tokens [zhou2024llm, liang2024can]. During summarization, we focus on items with low scores. Although best practices [hattie2007power] for feedback generally recommend including positive comments to support user motivation, analysis reports often require rapid actions. Therefore, feedback comments in this evaluation are restricted to checklist items that are less than or equal to “3" with the following prompt: “Based on the following evaluation results, generate feedback comments for improving the report in no more than 500 tokens. Preserve as much information from the original evaluation as possible, describe the issues concretely, and do not add any new information.”
6.1.2 Results
Table 6: Human and LLM ratings (1–5) for LLM-based evaluation results.
User Feedback Aspect Human Evaluation LLM Evaluation
Method 1 MESSALA Method 1 MESSALA
Overall Usefulness 3.00 4.00 4.03 4.35
Accuracy and Validity 4.00 3.67 4.15 4.41
Specificity and Concreteness 2.67 4.67 4.03 4.51
Support for Novices 4.00 4.67 3.48 3.65
Support for Veterans 3.00 3.67 4.08 4.35
The results are shown in Table 6. We also present examples of the feedback generated by the LLM in Appendix 0.E. Overall, MESSALA obtains higher ratings, except for “Accuracy and Validity". We describe the reason for each metric below.
Overall Usefulness: MESSALA was rated more useful than Method 1 because it more clearly pointed out concrete improvement points in analysis reports. As one participant noted, “Compared to Method 1, evaluation and suggestions for improvement by MESSALA are more specific and easier to understand.” – UserID 1. In contrast, feedback by Method 1 was readable but often inconsistent and generic: “The report is concise and approachable, but its feedback lacks consistency compared to MESSALA, reducing its usefulness.” – UserID 5.
Accuracy and Validity: Method 1 received a slightly higher score with 4.0 than MESSALA with 3.67. Although its feedback remained superficial, it was generally reasonable with few errors or contradictions, which contributed to favorable judgments. Whereas MESSALA is more detailed, it sometimes focuses on issues that the participants consider minor, leading to slightly lower scores: “MESSALA sometimes overemphasizes minor issues, resulting in overly detailed feedback on less important points.” – UserID 5. Overall, MESSALA can provide detailed evaluations but still struggles to prioritize the most critical issues from the perspective of human analysts. More fine-grained guidelines for each checklist item may improve this point.
Specificity and Concreteness: MESSALA outperformed Method 1 on this metric. The participants evaluated how feedback comments aligned each key point with its suggestion, making LLM outputs easier to translate into actionable items: “In MESSALA, key points and suggested improvements are mapped one-to-one, making action items easy to derive.” – UserID 5. Method 1 was reasonable but often lacked explicit links between evaluations and feedback, and its wording was sometimes vague. The participants preferred more direct expressions such as “missing,” “unclear,” or “insufficient”: “More direct feedback – such as ‘missing,’ ‘unclear,’ or ‘insufficient’ – is necessary rather than vague comments.” – UserID 2.
Support for Novices and Veterans: MESSALA was also rated more than Method 1 as supporting both novice and veteran analysts. Its feedback is structured for key points, and hence novice analysts can quickly identify which parts of a report should be revised: “Because MESSALA feedback is focused and concise, it’s more accessible to novices than Method 1, which requires reading the full text to understand.” – UserID 2. For veteran analysts, one-by-one feedback from MESSALA made it easier to extract relevant information than Method 1. However, the participant noted that both methods tend to misinterpret technical terminology and rely on indirect expressions, highlighting the need for stronger domain expertise in LLMs: “Both methods sometimes independently interpret technical terms, which obscures their intended meaning, and although they extract implicit information, their feedback is less direct and targeted than that of human experts when it comes to improving report quality.” -— UserID 5
6.2 Evaluation on Defect-injected Analysis Reports
This section conducts evaluation on defect-injected analysis reports to examine how each method identifies and explains defects and their reasons, following prior works [liu2023reviewergpt, d2024marg]. For effective evaluations of analysis reports, a method must accurately identify concrete points to be improved in the reports in order to refine the reports. Prior work [liang2024can] has pointed out a limitation that evaluations by LLMs are sometimes overly lenient and may fail to highlight critical defects. Our goal is thus to investigate whether MESSALA can mitigate this limitation and provide more actionable and meaningful feedback.
6.2.1 Defect-injected Analysis Reports
The defect-injected analysis report dataset is designed to evaluate whether an LLM can accurately identify defects in analysis reports. It consists of reports with realistic defects that are observed in SOC environments: for instance, missing contextual information, insufficient justification for judgments, flawed causal reasoning, and unclear or impractical response descriptions. The dataset also defines defect categories with their reference review comments. These reference review comments are used to assess whether the LLM can generate appropriate feedback.
The injected defects are explicitly labeled, and therefore, evaluations of the reports can reproducibly be conducted. Here, we define a defect as at least one of the 11 evaluation perspectives derived in Section 3 is missing or unsatisfied. Based on defect categories in prior works [liu2023reviewergpt, d2024marg] and the reference review comments collected from SOCs of the authors’ organization, we consolidated the original 11 evaluation perspectives into 4 defect categories with confirmations by SOC practitioners. Table 7 shows the consolidation between the 4 defect categories and the 11 perspectives.
Table 7: Defect categories and their corresponding practitioner-oriented evaluation perspectives
Defect Category Short Description Related Evaluation Perspectives
Opaque Decision Rationale Key judgments and conclusions lack sufficient justification or coherent reasoning, undermining confidence in the final assessment. (1), (10), (4)
Unverifiable or One-Sided Analysis Analytical claims are made without sufficient verification, such as baselines, comparisons, or alternative explanations. (8), (10), (4)
Context-Agnostic Technical Interpretation Technical observations are interpreted without adequate consideration of on-site context, such as asset roles, operational background, or constraints. (5), (6), (7), (9), (4)
Non-Actionable Outcome Presentation Analysis results fail to translate into concrete, prioritized actions or clearly articulated impact and escalation implications. (2), (3), (11), (4)
In addition, following these categories, we inject defects into the reports, and it is desirable for the LLM to generate feedback that captures the reference review comments corresponding to the defects. Such reference review comments are derived from real analysis reports, for example: “The report does not explain why the communication occurred, nor does it assess whether the communication or the involved endpoint is anomalous.” We selected a subset of analysis reports of high quality from standpoints of both human analysts and an LLM to inject defects. Each report was then modified according to predefined rules to inject one or more defects from the defect categories described above. During the above injections, only the relevant parts of each analysis report were modified, while the overall structure and writing style were preserved. These datasets were constructed using LLMs, whereby their outputs were reviewed and refined by the authors and the defect samples were further confirmed through reviews by SOC practitioners. We finally utilized 44 analysis reports as the dataset.
We conduct a preliminary validation to verify that quality deterioration of analysis reports caused by the injected defects is measurable. For each defect category, we compare evaluation scores between clean and defect-injected analysis reports derived from the same original report, and then identify whether score deteriorates in the corresponding evaluation perspectives while scores on unrelated evaluation perspectives remain unchanged. Defect injection and validation are conducted independently and the LLM performs scoring without access to defect labels or human verification results.
6.2.2 Experimental Setting
We use the defect-injected analysis report dataset described above to examine how each method identifies defects in analysis reports. We focus on whether it can generate feedback beyond superficial detection of defects and capture issues that should be improved according to the reference review comments.
The evaluation targets the feedback comments generated by the LLM for each defect-injected analysis report. We assess, through manual inspection, that these comments capture the reference review comments. Specifically, each feedback comment is categorized into three levels: (1) clearly covered, where it sufficiently captures the reference review comment; (2) partially covered, where it partially captures an essential part of the reference review comments despite differences in focus and wording; and (3) not covered, where it captures nothing of the reference review comment. The criteria for determining whether a comment covers defects are agreed upon by the authors in advance. As a baseline, we use Method 1 described in Section 5. Comparison with this baseline enables us to assess how MESSALA generates more meaningful feedback.
Table 8: Coverage results (left) and defect detection breakdown (right).
(a) Coverage of Reference Review Comments
Method Clearly Partially Not
MESSALA 29 9 6
Method1 16 15 13
(b) MESSALA: Defect Category Breakdown
Category
Partially
/ Clearly
Not
covered
Opaque Decision Rationale 10 2
Unverifiable / One-Sided 7 2
Context-Agnostic 12 2
Non-Actionable Outcome 9 0
6.2.3 Results
Table 8 presents the defect identification results for all methods. MESSALA identifies 86.4% of the injected defects, exceeding Method 1 across all defect categories. This performance difference reflects systematic discrepancies in how each method evaluates the relationship between conclusions, supporting evidence, contextual information, and decision-relevant actions.
Opaque Decision Rationale
This category captures whether an evaluation requires explicit and traceable reasoning from observations to conclusions, rather than accepting conclusions as sufficient by themselves.
Method 1 frequently assigns high scores to reports that explicitly state conclusions, even when the underlying causes or technical justifications are not explained. In Fig. 4, Method 1 evaluates a report as sufficient because it mentions device roles and associated vulnerabilities, despite the absence of explanations linking these factors to concrete operational impact.
In contrast, MESSALA assigns lower scores when such traceability is missing. As shown in Fig. 4, MESSALA identifies that the report does not explain how anomalies or state changes in the identified devices would affect building operations, even though technically correct descriptions are present. This behavior reflects MESSALA’s requirement that conclusions be supported by explicit evidence and interpretable reasoning.
Unverifiable or One-Sided Analysis
This category distinguishes whether evaluation favors plausible narratives or requires verifiable and reproducible analytical processes.
Method 1 tends to accept explanations that are internally consistent and operationally plausible, even when the analytical basis for those explanations is not documented. In the example shown in Fig.5, Method 1 accepts a false-positive determination based on high-level operational explanations without examining why the communication occurred.
In contrast, MESSALA assigns lower scores by explicitly questioning missing analyses of communication causality and endpoint behavior. As illustrated in Fig. 5, MESSALA points out that the report fails to justify how the observed traffic patterns correspond to legitimate business activity, thereby treating unverifiable reasoning as a defect rather than acceptable plausibility.
Figure 4: Comparison of Evaluation Comments: Opaque Decision Rationale category
Method 1 The report indicates that the device with IP address XXX is likely a maintenance laptop brought on-site, and explicitly requests confirmation as to whether it is a legitimate maintenance device. In addition, the report includes information on vulnerabilities related to the industrial software WebAccess Node running on the BA servers, providing a partial explanation of the devices’ criticality and associated risks.
MESSALA - Check Point X. Clarity of Device Role and Criticality: … - Check Point Y. Impact Description Based on Device Criticality: In the Potential Impact on the Building section, the report mentions the existence of vulnerabilities and the possibility of cyberattacks if the devices are not legitimate. However, it provides limited explanation of how anomalies or state changes in these devices would concretely affect building operations. The impact assessment remains at a high-level risk indication and lacks detailed descriptions of operational consequences.
Figure 5: Comparison of Evaluation Comments: Unverifiable or One-Sided Analysis category
Method 1 The report explains why the detected alert is considered a false positive and why no response is required, based on on-site operational context such as whether the affected products are in use and whether the destination IP addresses are used for business purposes. It assumes the validity of these high-level operational explanations without questioning why the communication occurred or whether the communication itself or the involved endpoint exhibits anomalous behavior. It also describes a concrete mitigation action, namely whitelisting, demonstrating an explanation aligned with actual operational practices.
MESSALA - Check Point X.Verification of Concrete Operational Context: … - Check Point Y. Insufficient Justification and Accountability in False-Positive Determination: The alert is concluded to be a false positive based on high-level statements such as “this poses no issue” or “the affected product is not in use,” yet these claims are not grounded in a concrete explanation of the actual operational context or system architecture. Moreover, although the IDS detected anomalous communication patterns (e.g., crafted two-byte characters), the report fails to explain how such patterns correspond to—or deviate from—legitimate business communications. Instead, it merely asserts that the traffic was “confirmed to be used for business purposes,” without providing sufficient technical or contextual justification to support the false-positive determination.
Similar phenomena were also observed for the Context-Agnostic Technical Interpretation and Non-Actionable Outcome Presentation categories (Table LABEL:tab:breakcate). Method 1 often assigns high scores to reports when technical descriptions are accurate or comprehensive, even if their feedback comments contain neither sufficient context nor actionable items. In contrast, MESSALA consistently assigns low scores to reports that are interpreted regardless of their contexts, even when explanations are technically correct, as well as comments without actionable items for decision-making.
6.3 Answers to RQ3
Our results show that MESSALA can qualitatively evaluate analysis reports and generate feedback comments aligned with human experts based on the evaluation criteria. In the multi-metric rating, feedback comments generated by MESSALA were consistently rated higher by human experts across all the categories except for Accuracy and Validity, indicating that the LLM can approximate judgment from the perspectives of human analysts as actionable and meaningful feedback comments.
7 Limitations and Ethical Consideration
7.0.1 Limitations
Our study has several limitations that need to be addressed in future work. First, we utilized private datasets for analysis reports. While pseudo analysis reports are also used to support reproducibility, the number of reports and their length are limited, and hence, we need to extend them into a more diverse setting. Second, the Analyst-wise checklist contains a large number of items, and the important evaluation items may differ depending on the reports. Future work will concentrate on a mechanism that automatically selects these items, although they were manually selected in this paper. Second, the number of SOC practitioners involved in the human gold evaluation was limited. To confirm the impact of MESSALA on reducing analysis workload and improving the quality of analysis reports in real-world environments, we also need to conduct further investigation with a larger number of participants. Finally, the evaluation perspectives in this paper, including the guidelines and defect categories, are based on the literature review and semi-structured interviews with practitioners and do not fully cover the evaluation processes and defect patterns across all real-world SOCs. We are in the process of refining feedback from versatile operation contexts.
7.0.2 Ethical Considerations
We discuss ethics in interviews and analysis reports below.
Interview: Regarding interview data in this study, we informed the participants in advance of the research objectives, interview procedures, scope of data use, and privacy protection policy, and obtained their consent. The participants were volunteers, and we explained that they could withdraw at anytime without any disadvantage under the above informed consent. Audio recordings of the interviews were used solely in transcription and qualitative analysis for the informed research purposes and were stored in a local environment accessible only to the authors. During the coding process, we removed any information that leads to re-identification, such as personal and organization names including their project details, and anonymized them only using pseudonym IDs.
Analysis Reports: Before using the analysis reports in this study, we explained the following points regarding confidentiality and research ethics to the stakeholders responsible for the SOC and obtained consent to conduct the experiments. In accordance with the principle of data minimization, two of the authors manually anonymized sensitive information, such as proper names, and examined the risk that individual devices and organizations could be re-identified. Only the processed data was then submitted to an LLM service for which our affiliation has a formal contract (i.e., prohibiting secondary use, third-party provision, and data retention) and obtained the security approvals. Access to the data and any use beyond the scope of this study were prohibited even within the authors’ affiliation, thereby reducing the risk arising from the experiments.
8 Conclusion
In this paper, we discussed evaluations of analysis reports in SOCs using LLMs. To this end, we first constructed the Analyst-wise checklist by identifying evaluation criteria for analysis reports in SOCs from multiple perspectives of human experts as the answer to RQ1 through literature review and interviews with SOC practitioners. Next, we designed MESSALA by further introducing the Granularization Guideline and the multi-perspective evaluation LLM in addition to the Analyst-wise checklist. MESSALA can maximize evaluations of analysis reports with feedback comments from the standpoint of human SOC practitioners as the answer to RQ2. We also conducted extensive experiments for quantitative evaluations of analysis reports with MESSALA, and then demonstrated that MESSALA can provide the closest evaluation results to those of the SOC practitioners compared with the baseline methods. Finally, we conducted another experiments for qualitative evaluations of analysis reports in terms of the feedback comments using LLMs to the reports. We then showed that MESSALA can qualitatively evaluate analysis reports and generate feedback comments aligned with human analysts based on the evaluation criteria as the answer to RQ3. We plan to conduct further experiments with diverse setting, including a larger number of samples and feedback comments from versatile operation contexts, as well as deployments on real-world SOC environments.
Appendix
Appendix 0.A Question Sheet for User Study and Qualitative Analysis
We utilized the question sheets shown in Table 9 and Table 10.
Table 9: Question Sheet for User Background and Preliminary User Study
Category No. Summary
User Information 1-1. How many years of experience do you have in security alert analysis?
1-2. What areas of security monitoring do you do? (IT (various industries), IoT (factories, electric power, cars, etc.))
1-3. What is the role of the SOC?
1-4. What is your specialty? (Cybersecurity, Software Engineering, Computer Science, etc.)
1-5. What types of security alerts do you primarily cover?
1-6. What types of reports do you provide to your customers?
1-7. How often do you report to your customers?
1-8. What is the process for creating a security alert analysis report?
1-9. Where do you get the information you need to create reports? (e.g., SIEMs, log management systems, manual investigations)
1-10. How long does it take to create a single report?
Preliminary User Study 2-1. Are there any tools or automation systems you use to make report creation more efficient? If so, what are they?
2-2. Are there any dissatisfactions or challenges with the tools you use when creating reports? If so, what are they?
2-3. Is there anything that would be useful to have tool support for in the near future, even if not fully automated?
2-4. Is there anything you think can be automated in reporting? Conversely, what are the parts that are difficult to automate?
2-5. Do you have high expectations for using LLMs when creating reports? What do you think about using LLMs for reporting?
2-6. Have you worked to establish common report formats and templates? If not, what prevents unifying the format?
2-7. Although report items are somewhat fixed, does the detailed content vary significantly depending on the event? Is it difficult to describe details due to reliance on experience?
2-8. Is there sufficient feedback on report content from customers or veteran analysts? If so, what kinds of comments have you received?
2-9. When reviewing reports created by others, what criteria do you use to judge their quality?
2-10. What do you think are the main challenges in evaluating analysis reports?
2-11. What do you think is the most important part of an analysis report?
2-12. What kinds of report items make decision-making easier?
2-13. Do you feel that the reports you create are sufficient compared to what they ideally should be? If not, what are the main reasons (e.g., time constraints, lack of information)?
2-14. Is contextual information from the field important in report content? If so, why, and what challenges exist in handling such information?
2-15. What kinds of information make report creation difficult, either due to their content or the process of collecting and organizing them?
2-16. What aspects of reporting do you find particularly challenging?
Table 10: User Feedback Evaluation Items for LLM-Generated Reports
Category No. Summary
Overall Usefulness F-1. Do you find the LLM-generated feedback useful overall for analysts?
Accuracy and Validity F-2. Does the feedback accurately identify errors or missing information in the report? How well does it align with expert-level feedback?
Specificity and Concreteness F-3. Does the evaluation provide concrete improvement points and reproducible advice?
Support for Novices F-4. Is the feedback understandable and helpful even for novice analysts?
Support for Veterans F-5. Is the feedback still useful for highly veteran analysts?
Appendix 0.B Analyst-wise checklist Item
We utilized the Checklist Items shown in Table 11 and Fig. 12.
Table 11: Analyst-wise Checklist Item Type Sample. This checklist summarizes essential items that should be confirmed in a typical security incident analysis report. An additional column labeled “Optional” indicates whether the item is conditionally required depending on the incident context—for example, some items may be primarily addressed by a dedicated Security Incident Response Team (SIRT).
Report Content Category Checklist Item Type Details Optional
Technical understanding and root cause clarification of the event Basic Alert Information Basic alert metadata such as timestamp, alert name, alert ID, and the detecting system that triggered the event. No
Event Description & Interpretation A clear description and interpretation of the observed event, including what happened, how it was detected, and why it is considered relevant. No
Root Cause Identification and explanation of the underlying technical or human causes that led to the incident. No
Incident Source The origin of the incident, such as the attack source, initiating asset, user, or external entity involved. No
Impacted Systems The systems, assets, services, or environments affected by the incident or potentially impacted. No
Communication Details Details of suspicious or relevant communications, including endpoints, protocols, destinations, content characteristics, and frequency. No
Vulnerability Information Information on known or suspected vulnerabilities related to the incident, if applicable. No
Decision support and action planning Impact Assessment An assessment of the operational impact, severity, business relevance, and associated risk of the incident. No
Confirmation Requests Specific items, assumptions, or questions that require confirmation from on-site staff, system owners, or other relevant teams. No
Response Actions Actions taken or recommended in response to the incident, including containment, mitigation, or monitoring measures. Yes
Recurrence Prevention & Lessons Learned Post-incident considerations such as preventive measures, control improvements, and lessons learned to avoid recurrence. Yes
Accountability and quality assurance of the investigation and analysis process Investigation & Analysis Methods The methods, tools, data sources, and analytical procedures used during the investigation and analysis of the event. No
Analysis Status The current analysis status of the incident, including whether it is ongoing, unresolved, or concluded, and the rationale for that status. No
Evidence & Supporting Data Preservation and documentation of supporting evidence, such as logs, alerts, configuration data, or artifacts used in the analysis. Yes
Related Policies, Guidelines & Standards Relevant internal policies, operational guidelines, or external standards applicable to the incident or its handling. Yes
Table 12: Analyst-wise Checklist Items. Each checklist item specifies a concrete review question under a checklist item type.
Checklist Item Type Checklist Item
Basic Alert Information Does the report include basic information about the alert under analysis, such as its content, time of occurrence, and location?
Analysis Status Are the current progress of the analysis and the next steps described?
Has it been confirmed whether the observed event is still ongoing and whether any related events are occurring?
Event Description & Interpretation Does the report draw a concrete conclusion based on the analysis of the observed event?
Are the key messages of the report and the rationale for its conclusions sufficiently explained?
If the supporting information is insufficient, does the report indicate that appropriate inquiries have been made to the recipient side?
Is the event interpreted and explained in line with the on-site operational context?
Investigation & Analysis Methods Does the report describe the methods used to assess the likelihood of an attack and its potential impact for the observed event?
Has the necessary analysis been conducted from multiple perspectives?
Are the analysis process and the decision criteria explained?
Has the event been checked by comparing its communications and device behavior with usual traffic and behavior patterns?
Does the report verify whether the observed behavior actually matches the attack pattern that the alert is designed to detect?
Have past alerts been investigated as well?
In addition to the alert in question, have related alerts in the surrounding time period been examined?
Has it been checked whether there are any anomalies in the temporal sequence of communications and behaviors before and after the alert-triggering event?
Based on the communications and behaviors associated with the alert, does the report conduct a detailed investigation and risk assessment to determine whether any abnormal actions succeeded and to assess the likelihood of an attack?
Has it been verified whether the observed actions or communications deviate from the expected role and normal operation of the device or user?
Have external reputation and vulnerability information sources been consulted, and has their consistency with the observed event been evaluated?
Has the event been evaluated in light of site-specific allowed operations and prohibited (“NG”) conditions?
Root Cause Has the root cause of the alert been identified and documented?
Does the report explain the analysis that led to the identification of the cause?
Is it distinguished whether the cause is human (manual operation, maintenance) or system-related (automatic/periodic processing, failure, malfunction, etc.)?
If the cause may be an attack, is the type of attack described concretely?
Incident Source Is detailed information about the source of the alert (e.g., the device’s on-site name and role) documented?
Has the importance of the source (device/user) been assessed?
Has the degree of suspiciousness of the source been evaluated?
Does the report describe any changes in the state of the source?
Does the report describe the current status of the source?
Has it been checked whether the communications performed by the device or user are appropriate given its intended role?
If the device or user is unknown, does the report provide a reasonable hypothesis about its possible identity or role?
Impacted Systems Is information about the alert’s communication destination documented in sufficient detail and without omissions?
Has the importance (criticality) of the destination resource been assessed?
Does the report describe any changes in the state of the communication destination?
Does the report describe the current status of the communication destination?
If the destination is an unknown device, does the report provide a reasonable hypothesis about its identity or role?
Communication Details Is the content of the alert-related communication documented in detail, including responses and whether each communication succeeded or failed?
Does the report explain how the communication relates to on-site operations?
Is the abnormality of the communication and the likelihood of an attack described and evaluated concretely?
Does the report include information about the network segments (e.g., VLAN IDs, segment names, and their respective roles)?
Impact Assessment Does the report concretely describe the impact of the event on the site/operational environment?
Does the report concretely describe the potential impact if the event were an actual attack?
Are the criteria and rationale used for the risk assessment explicitly stated?
Confirmation Requests If confirmation by the recipient is required, is the necessity of such confirmation and its rationale sufficiently explained?
Are specific items or questions to be confirmed by the recipient regarding this event described?
Are the requested confirmations or investigations realistically feasible for the recipient side?
Response Actions Is it clearly stated whether response actions are required, and is the rationale for this necessity sufficiently explained?
Are specific response actions for the current event described?
Does the report describe the expected effects of the proposed response actions and how those effects will be measured?
Recurrence Prevention & Lessons Learned Are concrete recurrence-prevention or improvement measures described?
Does the report include the expected effects of these preventive/improvement measures and how they will be evaluated or monitored?
Are the proposed recurrence-prevention or improvement measures realistic and actionable?
Does the report articulate lessons learned from the event and future response policies or courses of action?
Evidence & Supporting Data Is evidence supporting each post-incident measure clearly indicated?
Are the methods and sources for obtaining the evidence described?
Vulnerability Information Has known vulnerability information relevant to the event been collected?
Is the collected vulnerability information linked to and explained in relation to the analysis findings?
Does the report explain whether and how the identified vulnerabilities may apply to the observed event?
Related Policies, Guidelines & Standards Does the report identify any policies, guidelines, or industry standards that may have been violated?
Has the risk and impact of such non-compliance been assessed?
Appendix 0.C Granularization Guideline
We utilized the Guideline prompts shown in Fig. 6.
Figure 6: Overview of Prompts by Category in the Granularization Guideline.
Appendix 0.D Sample Prompt Used in MESSALA
We utilized the prompts shown in Fig. 7.
Figure 7: Overview of Our Prompt for Granularization and Evaluation. The red text indicates the distinctive characteristics of each prompt. Note that this prompt sample is constructed using general terminology, without referring to specific concepts such as the granularization guideline or Analyst-wise checklist.
Appendix 0.E LLM Evaluation Output Sample
This section presents the 5-point evaluation results and their corresponding reasons generated by each method (see Fig.8).
Figure 8: Examples of LLM-generated feedback for each method, generated by GPT-4o, shown alongside the corresponding sanitized report and checklist item.
Appendix 0.F Pseudo Analysis Report and Defect-injected Analysis Report Generation
This section presents example prompts used to generate the pseudo analysis reports in this paper, along with representative examples of the resulting pseudo analysis reports. It also includes sample defect-injected analysis reports.
Figure 9: Illustration of pseudo analysis report generation. The left image shows an example prompt used to generate a pseudo analysis report, while the right image presents the resulting pseudo analysis report produced by the LLM based on the prompt.
Figure 10: Comparison of the report excerpts before and after defect injection. Red text highlights the injected differences.
0.F.1 Defect-injected Rule
Below, we describe the defect injection rules for the Opaque Decision Rationale category. Similar rules are defined for the other categories, and defects are primarily injected into the reports through manual modification.
1) Preserve the final analytical conclusions: Maintain the original conclusions of the analysis report, such as the presence or absence of an attack, the overall risk assessment, and the final judgment outcome, without modification.
2) Remove assumptions and evaluation criteria underlying the decision: Intentionally eliminate or obscure the assumptions, comparison baselines, and evaluation criteria that support the analyst’s judgment, thereby concealing the decision-making framework.
3) Omit or weaken the logical linkage between evidence and conclusions: Remove explicit explanations that connect observed evidence, such as logs or communication behaviors, to the stated conclusions, creating gaps in the reasoning process.
4) Eliminate explicit causal expressions: Delete or replace phrases that clearly indicate causal reasoning, for example, “because,” “for the following reasons,” or “based on,” with abstract or non-specific expressions.
5) Abstract detailed analytical reasoning into vague summaries: Simplify concrete interpretations grounded in specific observations, such as correlations among multiple communication patterns, into high-level statements that merely state the inferred outcome.
6) Present judgments as assertive statements without justification: Express conclusions in a confident and definitive manner while providing insufficient or no supporting rationale.
7) Finalize a superficially complete but unverifiable report: Produce a report that appears complete in terms of conclusions but lacks sufficient reasoning for reviewers to assess the validity, soundness, or reproducibility of the analysis.